OpenStudy (anonymous):

Solve and find the solution set of: II x-2 I + 4 = 8 > note: this is about EQUATIONS AND INEQUALITIES INVOLVING THE ABSOLUTE VALUE

OpenStudy (anonymous):

when x is +ve, x-2+4=8 x+2=8 x=6 when x is -ve, -x+2+4=8 -x+6=8 -x=2 x=-2

OpenStudy (anonymous):

so x = 6 x = -2

OpenStudy (anonymous):

or could it be x=-10 x=14 ???

OpenStudy (anonymous):

i'm confused...

OpenStudy (saranya):

I x-2 I + 4 = 8 | x-2 | = 4 x = 6 (when +ve) I x-2 I + 4 = 8 | x-2 | = 4 x =- 2 (when -ve)

OpenStudy (saranya):

why yu r confused ?

OpenStudy (anonymous):

because there are two bars

OpenStudy (anonymous):

oh yeah.. there's a bar after 4 --- i forgot to add that so the given is now: II x-2 I + 4 I = 8

OpenStudy (anonymous):

Ix-2I+4>0

OpenStudy (anonymous):

what is the result?

OpenStudy (maheshmeghwal9):

\[\color{red}{|x-2|+4=8.}\] \[\color{blue}{case-I}\]\[\color{purple}{x-2}+4=8.\]\[\implies x=8-4+2=6.\] \[\color{blue}{case-II}\]\[\color{purple}{-(x-2)}+4=8.\]\[\implies -x+2+4=8 \implies -x=8-4-2=2.\]\[\implies -x=2 \implies x=-2.\]So; \[\color{green}{x=6\space and -2.}\]

OpenStudy (maheshmeghwal9):

Sol. for ur first question second one is also easy just concentrate on concept that |a| = a ; if a>0 -a ; if a<0

OpenStudy (maheshmeghwal9):

For ur actual question; case- I +[(x-2)+4]=8 case-II +[-(x-2)+4]=8. case-III -[(x-2)+4]=8. case-IV -[-(x-2)+4]=8. \[\huge{\text{Solve these on ur own.}}\]

OpenStudy (maheshmeghwal9):

there will be 4 solutions.

OpenStudy (maheshmeghwal9):

Any doubt?

OpenStudy (maheshmeghwal9):

If yes then tell actual doubt:)